# Joint Pruning and Channel-wise Mixed-Precision Quantization for Efficient Deep Neural Networks

### v1

@@ 5.4 Deployment
In the leftmost plot of Figure 6, we present the accuracy vs. execution cycles trade-off of each model, when deployed on the MPIC hardware. The rightmost plot presents instead the results obtained by deploying the trained models on the NE16 accelerator. We also report the results of models obtained with the “wrong” cost model for both targets, to show the importance of hardware-awareness. Indeed, while using a cost model tailored to a different hardware does not impact the results significantly in the MPIC case, mainly thanks to the flexibility of this CPU-based platform, the benefit of matching the complexity metric to the actual hardware platform are significant in the case of NE16. For example, at iso-accuracy with the baseline w2a8, we obtain a model with a 22.14% increase and a 50.43% decrease in number of cycles on the NE16 hardware when using MPIC and NE16 cost models, respectively. This is due to the more complex dataflow and spacial parallelism of this accelerator, which make it more convenient to assign the same bit-width to entire chunks of channels, otherwise some computational bandwidth is wasted and the latency gain is reduced. This is taken into account by the NE16 model, but not by the MPIC one, thus leading to suboptimal solutions.

### v2

@@ 5.4 Deployment
In the leftmost plot of Fig. 6, we present the accuracy vs. execution cycles trade-off of each model, when deployed on the MPIC hardware. The rightmost plot presents instead the results obtained by deploying the trained models on the NE16 accelerator. We also report the results of models obtained with the “wrong” cost model for both targets, to show the importance of hardware-awareness. Indeed, while using a cost model tailored to a different hardware does not impact the results significantly in the MPIC case, mainly thanks to the flexibility of this CPU-based platform, the benefit of matching the complexity metric to the actual hardware platform are significant in the case of NE16. For example, at iso-accuracy with the baseline w2a8, we obtain a model with a 22.14% increase and a 50.43% decrease in number of cycles on the NE16 hardware when using MPIC and NE16 cost models, respectively. This is due to the more complex dataflow and spacial parallelism of this accelerator, which make it more convenient to assign the same bit-width to entire chunks of channels, otherwise some computational bandwidth is wasted and the latency gain is reduced. This is taken into account by the NE16 model, but not by the MPIC one, thus leading to suboptimal solutions.